OpenStudy (staceyg):

What's 75% of something worth 38% .

OpenStudy (chosenmatt):

welcome to openstudy =)

OpenStudy (anonymous):

welcome to openstudy :)

OpenStudy (staceyg):

@chosenmatt @shifuyanli Can you guys answer a different question?

OpenStudy (anonymous):

yea sure

OpenStudy (chosenmatt):

sure

OpenStudy (anonymous):

make sure you post it in a new thread please :)

OpenStudy (staceyg):

State the property that justifies the statement. If 3x - 4 =10, then 3x = 14

OpenStudy (anonymous):

nvm

OpenStudy (anonymous):

what do you think?

OpenStudy (staceyg):

subtraction?

OpenStudy (anonymous):

nope :) not that kind of property

OpenStudy (anonymous):

I think its addition

OpenStudy (anonymous):

yes @lexie<3 but let @staceyg try it

OpenStudy (staceyg):

The choices are either subtraction, addition, multiplication, or division property of equality.

OpenStudy (staceyg):

How did you get that @lexie<3

OpenStudy (anonymous):

well what do you think it is if it is not subtraction?

OpenStudy (anonymous):

what is the opposite of subtraction?

OpenStudy (anonymous):

3x-4=10 3x=10+4 3x=14

OpenStudy (staceyg):

i know but how would you get that answer?

OpenStudy (anonymous):